OpenStudy (anonymous):

Can somebody explain me on how to prove whether this is linear transformation or not...with the working solutin of course :( .. L(x,y)=(x+y,x-y) ?

OpenStudy (anonymous):

Do you know what two properties you need to show?

OpenStudy (anonymous):

I don't know.Perhaps the only thing I know is the most basic formula which is L(u+v)=L(u)+L(v) and L(ku)= kL(u).Please help me..I'm poor in this..

OpenStudy (anonymous):

Good. Here, your (x,y) takes the place of u. So, start with L(a+b,c+d), and work with the definition to get L(a,c) + L(b,d). The other property is similar. So, give it a try? I'll point out if you're correct or if you make a mistake.